OpenStudy (anonymous):

What is the antiderivative of x(1-x)?

OpenStudy (freckles):

distribute first

OpenStudy (freckles):

That might make things easy to see

OpenStudy (freckles):

\[\int\limits_{}^{}x^n dx=\frac{x^{n+1}}{n+1}+C , n \neq -1 \]